# Navigation Around Unknown Space Objects Using Visible-Thermal Image Fusion

### v1

@@ 1 Introduction
This work focuses on uncooperative targets, which are not designed to interface with an inspector in any way. Uncooperative targets are considered “known” if a 3D model of the target’s shape is available to the inspector prior to rendezvous. An inspector can estimate its position and attitude relative to the target (known as pose estimation) by matching 2D monocular images to the 3D model [sharma_comparative_2016, pauly_survey_2023] or by matching a 3D point cloud obtained from lidar to the 3D model [ruel_target_2008, opromolla_lidar_2014, liu_point_2016, woods_lidar_2016, perfetto_lidar_2019]. Cassinis et al. provide a general survey of these spacecraft pose estimation techniques [cassinis_review_2019].
However, a target is not always “known” prior to the mission. These uncooperative and unknown targets are much more difficult to handle because a model of the target needs to be created while simultaneously being used to find the relative pose of the inspector. This class of algorithms is known as Simultaneous Localization and Mapping (SLAM) and is a prominent focus of research in robotics [slam_handbook_2025]. Lichter et al. [lichter_state_2004] first explored this problem in the context of space objects using collaborative 3D vision sensors (assumed to be uniformly distributed around the target), with a Kalman filter used to estimate the state, shape, and mass parameters of an unknown target. Sonnenburg et al. [sonnenburg_ekf_2010] used a set of stereo cameras to find the position of 3D features (high-contrast corners) on the target. These are inputs to an extended Kalman filter, which finds the relative pose of the inspector. A monocular camera can also be used to implement a similar system [augestein_improved_2011]. By tracking 2D features on the images of the target over multiple frames, the features can be triangulated and positioned in 3D space. This technique is known as structure from motion (SfM). Tweddle [tweddle_thesis_2013] implemented a SLAM algorithm assuming a stationary inspector that relies on a factor graph for mapping and pose estimation. A more in-depth description of this technique is found in Subsection 2.4. Setterfield et al. [setterfield_mapping_2018] expanded on Tweddle’s work so that it can handle a moving inspector. Orbital dynamics priors can also be incorporated into the SLAM estimate. Dor et al. [dor_astroslam_2024] formulated a SLAM system to navigate around an unknown small asteroid in orbit around the Sun using these priors. The effects of the Sun’s gravity, the asteroid’s gravity, and solar radiation pressure are used to improve the estimated trajectory of the inspector and the map of the asteroid. One significant problem with these inspection missions is the large operating ranges between the inspector and target. Such ranges create degeneracies in SfM algorithms caused by weak-perspective projection, center-pointing trajectories, and dominant planar geometry of target satellites. These problems have been investigated by Florez et al. [florez_initialization_2025], who created a monocular pipeline robust to these degeneracies for spacecraft inspection.
@@
Given these considerations, an additional sensor should be used by the inspector to improve navigation robustness. Lidar emits infrared lasers to measure the bearing and range of every point on the target. Unlike a camera, this active sensing method has the advantage of directly measuring the range to the target and is independent of illumination conditions. However, this also means that lidars use more power and are often significantly heavier and larger than a camera—constraints that are often difficult to accommodate in space vehicle design. Strong enough lidars also have the potential to damage fragile sensors on the target [bi_failure_2024].
@@
Imagery from TIR cameras has been shown to be useful on flight demonstrations to the ISS [ruel_space_2012, cavrois_liris_2015], as shown in Fig. 2. These missions demonstrate that using a TIR imager, the ISS is clearly visible over the entire orbit, even when fully eclipsed. The use of TIR has been studied for relative navigation around known and unknown RSOs, with a focus on how to utilize information from both VIS and TIR cameras simultaneously. Combining the information from VIS and TIR images creates a solution that has the advantages of each while hopefully avoiding their disadvantages. Civardi et al. [civardi_small_2021] proposed a method to combine the VIS and TIR information to navigate around an unknown asteroid. SLAM is run separately for each sensor, and then the estimated relative pose from each SLAM solution is combined in an extended Kalman filter. While this method combines the estimation from the VIS and TIR sensors, it does so in a way that decouples the information between the two sensors. In other words, the features detected in the VIS image are not improving the navigation estimate obtained from the TIR image and vice versa. An approach that has been investigated for pose estimation around a known object [mcbryde_spacecraft_2018, civardi_generation_2024, civardi_uncooperative_2024] is the fusion of the VIS and TIR images at the pixel level before they are input into any navigation algorithm. This approach uses the joint information in both sensors to create a navigation estimate, rather than fusing two decoupled pose estimates.
@@
Figure 2: The ISS as seen from the TIR imager of the TriDAR vision system during rendezvous operations on STS-131 [ruel_space_2012].
@@ 2.1 Relative Orbital Dynamics
The relative motion of the inspector around the target can be linearized to create a set of coupled ordinary differential equations, named the Hill-Clohessy-Wiltshire (HCW) equations [clohessy_terminal_1960]
@@
When solving the HCW equations, constraints can be imposed that bound the relative orbit of the inspector around the target, such that these orbits take the form of closed ellipses that repeat over time in the Hill frame [schaub_analytical_2018]. These solutions take the form of
@@
where $A_{0}$, $B_{0}$, $\alpha$, $\beta$, and $y_{\text{off}}$ are integration constants to be satisfied by initial conditions. Similar bounded relative orbits have also been shown to exist for cases where the target has an elliptical orbit around the parent body, as well as where a non-spherical gravitational potential exists (such as LEO) [schaub_analytical_2018].
@@ 2.2 Eclipse Conditions
A significant issue with visible cameras is their inability to see parts of the target that are not illuminated by the Sun. This occurs when the target is in between the inspector and the Sun, such that the target is backlit. However, a far more significant and common scenario is when Earth is eclipsing the target. To get a sense of how significant this effect is, we define the beta angle, $\beta$, as the angle between the solar vector and its projection onto the orbital plane [gilmore_spacecraft_2002]. This relationship is shown in Fig. 3.
@@ 2.3 Thermal Dynamics
In order to simulate a spacecraft in the thermal band, it is important to first understand the sources of thermal radiation in LEO [gilmore_spacecraft_2002]. The sources we consider are internal heating, direct solar, Earth albedo, and Earth IR.
@@ 2.4 Simultaneous Localization and Mapping
Simultaneous Localization and Mapping (SLAM) is the process of creating a map of an unknown environment while finding the position of an agent within that map. A detailed and modern report on the theory and methods can be found in the SLAM Handbook [slam_handbook_2025].
In this work, Kimera [kimera, kimera2] is used as the SLAM algorithm of choice. Kimera is a feature-based SLAM framework that distills priors, measurements, and dynamics into a factor graph to jointly optimize both the inspector’s relative motion and the 3D positions of detected feature points on the target. In monocular mode, all Kimera requires is monocular camera images and IMU measurements. Two modules within Kimera are used: Kimera-VIO and Kimera-RPGO.
@@ 2.5 VIS/TIR Fusion
Before image fusion, some pre-processing must take place. The input VIS and TIR images are often not the same resolution, with TIR usually a substantially lower resolution than VIS. However, the two images need to have matching resolution for the fusion methods to combine them. Civardi et al. [civardi_generation_2024] show that upscaling the TIR images to match the resolution of the VIS creates a fused image that is least sensitive to noise, as compared to downscaling the VIS images to match the TIR resolution. This approach is not only preferred because it is more robust to noise, but also because the higher resolution of the fused image allows for more precise placement of feature points, creating a more accurate navigation solution.
Civardi et al. also present an analysis of different image processing and fusion techniques to combine two separate VIS and TIR images into a fused image in the context of uncooperative spacecraft. Their analysis compares how closely the fused images retain the information from the input VIS and TIR images in the presence of noise, as well as the runtime of each algorithm. They find that two image fusion methods create the best resultant composites with a fast runtime: Anisotropic Diffusion-Based Fusion (ADF) [bavirisetti_fusion_2016] and Two-scale Image Fusion using Saliency Detection (TSIFSD) [bavirisetti_twoscale_2016]. We compare the performance of TSIFSD and ADF in the context of SLAM around an unknown spacecraft.
@@
2.5.1 Two-scale Image Fusion using Saliency Detection (TSIFSD)  [bavirisetti_twoscale_2016]
@@
2.5.2 Anisotropic Diffusion-Based Fusion (ADF) [bavirisetti_fusion_2016]
Anisotropic diffusion [perona_anisotropic_1990] is an iterative process that smooths an image at homogeneous regions, but preserves edges. This is used to separate the input VIS and TIR images into base layers and detail layers, similar to TSIFSD. The detail layers are then combined using a Karhunen–Loève transform [hotelling_analysis_1933, karhunen_lineare_1947, loeve_functions_1948], which decorrelates the information between the VIS and TIR layers into principal components. The strongest components are then combined to obtain the fused detail layer. After the base layers are combined with a simple average, the fused base and detail layers are added together.
@@ 3.2.2 Environment
In the VIS simulation, the two significant illumination sources are the Sun and Earth albedo. The Sun is simulated as an inertially fixed point source infinitely far away. The Earth is simulated photo-realistically as a sphere, with land, water, and cloud textures applied from satellite imagery. Atmospheric scattering is also included in this simulation. The rendering of Earth has a significant effect on the lighting of the target spacecraft, often filling in the harsh shadows left by the Sun. This work does not plan to study the effects on tracking performance of Earth in the background of images, so the direct visibility of Earth is excluded from the renderings. However, it has been shown that the use of TIR imagery is far more robust to the clutter of Earth in the background [fiengo_multispectral_2019].
@@ 3.2.3 Camera
Another significant camera artifact is noise. A small amount of Gaussian noise is added to the VIS camera to simulate the electronic noise inherent to the camera. Most TIR cameras use an uncooled microbolometer imager which, unlike other types of TIR imagers, do not require a cryogenic cooler. However, the lack of cooling creates significant noise in the form of fixed-pattern noise (FPN) consisting of vertical and horizontal streaks across the image [Barral_FPN_2024]. This FPN is added to the rendered TIR images, along with standard Gaussian noise.
@@ 3.4 Relative Dynamics Measurements
The dynamics of an on-orbit inspector behave very differently from the dynamics of a ground robot or UAV, however. Against Kimera’s assumptions, the environment being mapped (the target RSO in this case) may not have an attitude that is inertially fixed if it has some rotational velocity. Similarly, the target’s inertial position is certainly not fixed as it is orbiting Earth. As a result, the SLAM formulation needs to be adjusted to estimate the rotational velocity of the target and to incorporate the relative orbital dynamics between the target and inspector. Dor et al. formulate a SLAM framework that handles this complication of frames [dor_astroslam_2024].
In order to simplify the problems arising from a non-inertially-fixed target to instead focus on the problem of visible-thermal fusion, this work assumes we have some noisy translational acceleration and rotational velocity measurements relative to the body frame of the target. Rotational velocity measurements are simulated using the ground truth relative trajectory with added noise, random walk, and bias instability to match measurement errors seen from a gyroscope. Relative acceleration measurements are also simulated using the ground truth relative trajectory. Relative orbital dynamics are the primary method to find these acceleration measurements [dor_astroslam_2024]. As a result, errors are added to the simulated acceleration measurements to represent errors that would be found in the propagation of relative dynamics. Orbital propagation errors on the order of hundreds of meters over a propagation period of 24 hours are seen with common analytical propagation models [vallado_revisiting_2006]. The simulated relative acceleration measurements are therefore created with errors of this magnitude when integrated twice. MATLAB is used to generate the data using the parameters shown in Table 2.
@@ 4.1 Image Processing
After images are generated in Blender, preprocessing is applied to increase contrast and reduce noise in the images before they are fed to Kimera. For each VIS and TIR frame, a threshold mask is used to separate the target from the dark background to minimize amplification of noise. Contrast Limited Adaptive Histogram Equalization (CLAHE) [zuiderveld_clahe_1994] is then applied to each VIS and TIR frame to improve local contrasts over the image and enhance the definition of edges, a common technique used to improve feature detection performance.
@@ 5 Results
For each run, the position and rotation errors between the estimated and ground truth relative trajectories are found using evo [grupp_evo_2017], which averages them over the whole trajectory. These results are shown in Table 4. The maximum errors over each trajectory are also calculated and included in Table 5. Table 6 shows the number of loop closures detected out of the 1,661 frames in each run. The percent of frames during which successful tracking and mapping is achieved is shown in Table 7. The average number of inlier feature matches per frame is presented in Table 8, only including frames which had successful tracking and mapping.
@@ 5.3 Fused VIS/TIR
The results also indicate that TSIFSD generally performs better than ADF. TSIFSD retains more detail in the fused images, shown by the significantly higher number of feature matches per frame. Accompanied by the fact that TSIFSD has a runtime almost an order of magnitude less than ADF [bavirisetti_twoscale_2016], TSIFSD is the recommended method of fusion in this analysis.
@@ 6 Conclusion
While this work shows the practicality of visible-thermal image fusion, some assumptions are made here that should be replaced before the use of this algorithm in flight. Namely, the generation of relative translational acceleration and rotational velocity measurements should realistically be replaced with a system that creates real-time estimates of the inspector’s relative motion using orbital dynamics. A formulation created by Dor et al. [dor_astroslam_2024] can be used to remove the assumptions used here.
The navigation methods presented here rely on image-level fusion between visible and thermal-infrared. However, another method of fusion could combine the visible and thermal information in the back-end optimization of the SLAM algorithm. Features can be tracked on visible and thermal images separately, followed by the 3D mapping of each landmark in a joint visible-thermal map of the target. In this framework, the stereo baselines between the visible and thermal cameras can be used to calculate the depth of features that are common between the two cameras. This navigation framework has already been tested for terrestrial navigation [qin_BVT-SLAM_2024] and may offer an even better visible-thermal fusion method in the context of on-orbit inspection.

### v2

@@ 1 Introduction
This work focuses on uncooperative targets, which are not designed to interface with an inspector in any way. Uncooperative targets are considered “known” if a 3D model of the target’s shape is available to the inspector prior to rendezvous. An inspector can estimate its position and attitude relative to the target (known as pose estimation) by matching 2D monocular images to the 3D model [1, 2] or by matching a 3D point cloud obtained from lidar to the 3D model [3, 4, 5, 6, 7]. Cassinis et al. provide a general survey of these spacecraft pose estimation techniques [8].
However, a target is not always “known” prior to the mission. These uncooperative and unknown targets are much more difficult to handle because a model of the target needs to be created while simultaneously being used to find the relative pose of the inspector. This class of algorithms is known as Simultaneous Localization and Mapping (SLAM) and is a prominent focus of research in robotics [9]. Lichter et al. [10] first explored this problem in the context of space objects using collaborative 3D vision sensors (assumed to be uniformly distributed around the target), with a Kalman filter used to estimate the state, shape, and mass parameters of an unknown target. Sonnenburg et al. [11] used a set of stereo cameras to find the position of 3D features (high-contrast corners) on the target. These are inputs to an extended Kalman filter, which finds the relative pose of the inspector. A monocular camera can also be used to implement a similar system [12]. By tracking 2D features on the images of the target over multiple frames, the features can be triangulated and positioned in 3D space. This technique is known as structure from motion (SfM). Tweddle [13] implemented a SLAM algorithm assuming a stationary inspector that relies on a factor graph for mapping and pose estimation. A more in-depth description of this technique is found in Subsection 2.4. Setterfield et al. [14] expanded on Tweddle’s work so that it can handle a moving inspector. Orbital dynamics priors can also be incorporated into the SLAM estimate. Dor et al. [15] formulated a SLAM system to navigate around an unknown small asteroid in orbit around the Sun using these priors. The effects of the Sun’s gravity, the asteroid’s gravity, and solar radiation pressure are used to improve the estimated trajectory of the inspector and the map of the asteroid. One significant problem with these inspection missions is the large operating ranges between the inspector and target. Such ranges create degeneracies in SfM algorithms caused by weak-perspective projection, center-pointing trajectories, and dominant planar geometry of target satellites. These problems have been investigated by Florez et al. [16], who created a monocular pipeline robust to these degeneracies for spacecraft inspection.
@@
Given these considerations, an additional sensor should be used by the inspector to improve navigation robustness. Lidar emits infrared lasers to measure the bearing and range of every point on the target. Unlike a camera, this active sensing method has the advantage of directly measuring the range to the target and is independent of illumination conditions. However, this also means that lidars use more power and are often significantly heavier and larger than a camera—constraints that are often difficult to accommodate in space vehicle design. Strong enough lidars also have the potential to damage fragile sensors on the target [17].
@@
Imagery from TIR cameras has been shown to be useful on flight demonstrations to the ISS [18, 19], as shown in Fig. 2. These missions demonstrate that using a TIR imager, the ISS is clearly visible over the entire orbit, even when fully eclipsed. The use of TIR has been studied for relative navigation around known and unknown RSOs, with a focus on how to utilize information from both VIS and TIR cameras simultaneously. Combining the information from VIS and TIR images creates a solution that has the advantages of each while hopefully avoiding their disadvantages. Civardi et al. [20] proposed a method to combine the VIS and TIR information to navigate around an unknown asteroid. SLAM is run separately for each sensor, and then the estimated relative pose from each SLAM solution is combined in an extended Kalman filter. While this method combines the estimation from the VIS and TIR sensors, it does so in a way that decouples the information between the two sensors. In other words, the features detected in the VIS image are not improving the navigation estimate obtained from the TIR image and vice versa. An approach that has been investigated for pose estimation around a known object [21, 22, 23] is the fusion of the VIS and TIR images at the pixel level before they are input into any navigation algorithm. This approach uses the joint information in both sensors to create a navigation estimate, rather than fusing two decoupled pose estimates.
@@
Figure 2: The ISS as seen from the TIR imager of the TriDAR vision system during rendezvous operations on STS-131 [18].
@@ 2.1 Relative Orbital Dynamics
The relative motion of the inspector around the target can be linearized to create a set of coupled ordinary differential equations, named the Hill-Clohessy-Wiltshire (HCW) equations [24]
@@
When solving the HCW equations, constraints can be imposed that bound the relative orbit of the inspector around the target, such that these orbits take the form of closed ellipses that repeat over time in the Hill frame [25]. These solutions take the form of
@@
where $A_{0}$, $B_{0}$, $\alpha$, $\beta$, and $y_{\text{off}}$ are integration constants to be satisfied by initial conditions. Similar bounded relative orbits have also been shown to exist for cases where the target has an elliptical orbit around the parent body, as well as where a non-spherical gravitational potential exists (such as LEO) [25].
@@ 2.2 Eclipse Conditions
A significant issue with visible cameras is their inability to see parts of the target that are not illuminated by the Sun. This occurs when the target is in between the inspector and the Sun, such that the target is backlit. However, a far more significant and common scenario is when Earth is eclipsing the target. To get a sense of how significant this effect is, we define the beta angle, $\beta$, as the angle between the solar vector and its projection onto the orbital plane [26]. This relationship is shown in Fig. 3.
@@ 2.3 Thermal Dynamics
In order to simulate a spacecraft in the thermal band, it is important to first understand the sources of thermal radiation in LEO [26]. The sources we consider are internal heating, direct solar, Earth albedo, and Earth IR.
@@ 2.4 Simultaneous Localization and Mapping
Simultaneous Localization and Mapping (SLAM) is the process of creating a map of an unknown environment while finding the position of an agent within that map. A detailed and modern report on the theory and methods can be found in the SLAM Handbook [9].
In this work, Kimera [27, 28] is used as the SLAM algorithm of choice. Kimera is a feature-based SLAM framework that distills priors, measurements, and dynamics into a factor graph to jointly optimize both the inspector’s relative motion and the 3D positions of detected feature points on the target. In monocular mode, all Kimera requires is monocular camera images and IMU measurements. Two modules within Kimera are used: Kimera-VIO and Kimera-RPGO.
@@ 2.5 VIS/TIR Fusion
Before image fusion, some pre-processing must take place. The input VIS and TIR images are often not the same resolution, with TIR usually a substantially lower resolution than VIS. However, the two images need to have matching resolution for the fusion methods to combine them. Civardi et al. [22] show that upscaling the TIR images to match the resolution of the VIS creates a fused image that is least sensitive to noise, as compared to downscaling the VIS images to match the TIR resolution. This approach is not only preferred because it is more robust to noise, but also because the higher resolution of the fused image allows for more precise placement of feature points, creating a more accurate navigation solution.
Civardi et al. also present an analysis of different image processing and fusion techniques to combine two separate VIS and TIR images into a fused image in the context of uncooperative spacecraft. Their analysis compares how closely the fused images retain the information from the input VIS and TIR images in the presence of noise, as well as the runtime of each algorithm. They find that two image fusion methods create the best resultant composites with a fast runtime: Anisotropic Diffusion-Based Fusion (ADF) [29] and Two-scale Image Fusion using Saliency Detection (TSIFSD) [30]. We compare the performance of TSIFSD and ADF in the context of SLAM around an unknown spacecraft.
@@
2.5.1 Two-scale Image Fusion using Saliency Detection (TSIFSD)  [30]
@@
2.5.2 Anisotropic Diffusion-Based Fusion (ADF) [29]
Anisotropic diffusion [31] is an iterative process that smooths an image at homogeneous regions, but preserves edges. This is used to separate the input VIS and TIR images into base layers and detail layers, similar to TSIFSD. The detail layers are then combined using a Karhunen–Loève transform [32, 33, 34], which decorrelates the information between the VIS and TIR layers into principal components. The strongest components are then combined to obtain the fused detail layer. After the base layers are combined with a simple average, the fused base and detail layers are added together.
@@ 3.2.2 Environment
In the VIS simulation, the two significant illumination sources are the Sun and Earth albedo. The Sun is simulated as an inertially fixed point source infinitely far away. The Earth is simulated photo-realistically as a sphere, with land, water, and cloud textures applied from satellite imagery. Atmospheric scattering is also included in this simulation. The rendering of Earth has a significant effect on the lighting of the target spacecraft, often filling in the harsh shadows left by the Sun. This work does not plan to study the effects on tracking performance of Earth in the background of images, so the direct visibility of Earth is excluded from the renderings. However, it has been shown that the use of TIR imagery is far more robust to the clutter of Earth in the background [35].
@@ 3.2.3 Camera
Another significant camera artifact is noise. A small amount of Gaussian noise is added to the VIS camera to simulate the electronic noise inherent to the camera. Most TIR cameras use an uncooled microbolometer imager which, unlike other types of TIR imagers, do not require a cryogenic cooler. However, the lack of cooling creates significant noise in the form of fixed-pattern noise (FPN) consisting of vertical and horizontal streaks across the image [36]. This FPN is added to the rendered TIR images, along with standard Gaussian noise.
@@ 3.4 Relative Dynamics Measurements
The dynamics of an on-orbit inspector behave very differently from the dynamics of a ground robot or UAV, however. Against Kimera’s assumptions, the environment being mapped (the target RSO in this case) may not have an attitude that is inertially fixed if it has some rotational velocity. Similarly, the target’s inertial position is certainly not fixed as it is orbiting Earth. As a result, the SLAM formulation needs to be adjusted to estimate the rotational velocity of the target and to incorporate the relative orbital dynamics between the target and inspector. Dor et al. formulate a SLAM framework that handles this complication of frames [15].
In order to simplify the problems arising from a non-inertially-fixed target to instead focus on the problem of visible-thermal fusion, this work assumes we have some noisy translational acceleration and rotational velocity measurements relative to the body frame of the target. Rotational velocity measurements are simulated using the ground truth relative trajectory with added noise, random walk, and bias instability to match measurement errors seen from a gyroscope. Relative acceleration measurements are also simulated using the ground truth relative trajectory. Relative orbital dynamics are the primary method to find these acceleration measurements [15]. As a result, errors are added to the simulated acceleration measurements to represent errors that would be found in the propagation of relative dynamics. Orbital propagation errors on the order of hundreds of meters over a propagation period of 24 hours are seen with common analytical propagation models [37]. The simulated relative acceleration measurements are therefore created with errors of this magnitude when integrated twice. MATLAB is used to generate the data using the parameters shown in Table 2.
@@ 4.1 Image Processing
After images are generated in Blender, preprocessing is applied to increase contrast and reduce noise in the images before they are fed to Kimera. For each VIS and TIR frame, a threshold mask is used to separate the target from the dark background to minimize amplification of noise. Contrast Limited Adaptive Histogram Equalization (CLAHE) [38] is then applied to each VIS and TIR frame to improve local contrasts over the image and enhance the definition of edges, a common technique used to improve feature detection performance.
@@ 5 Results
For each run, the position and rotation errors between the estimated and ground truth relative trajectories are found using evo [39], which averages them over the whole trajectory. These results are shown in Table 4. The maximum errors over each trajectory are also calculated and included in Table 5. Table 6 shows the number of loop closures detected out of the 1,661 frames in each run. The percent of frames during which successful tracking and mapping is achieved is shown in Table 7. The average number of inlier feature matches per frame is presented in Table 8, only including frames which had successful tracking and mapping.
@@ 5.3 Fused VIS/TIR
The results also indicate that TSIFSD generally performs better than ADF. TSIFSD retains more detail in the fused images, shown by the significantly higher number of feature matches per frame. Accompanied by the fact that TSIFSD has a runtime almost an order of magnitude less than ADF [30], TSIFSD is the recommended method of fusion in this analysis.
@@ 6 Conclusion
While this work shows the practicality of visible-thermal image fusion, some assumptions are made here that should be replaced before the use of this algorithm in flight. Namely, the generation of relative translational acceleration and rotational velocity measurements should realistically be replaced with a system that creates real-time estimates of the inspector’s relative motion using orbital dynamics. A formulation created by Dor et al. [15] can be used to remove the assumptions used here.
The navigation methods presented here rely on image-level fusion between visible and thermal-infrared. However, another method of fusion could combine the visible and thermal information in the back-end optimization of the SLAM algorithm. Features can be tracked on visible and thermal images separately, followed by the 3D mapping of each landmark in a joint visible-thermal map of the target. In this framework, the stereo baselines between the visible and thermal cameras can be used to calculate the depth of features that are common between the two cameras. This navigation framework has already been tested for terrestrial navigation [40] and may offer an even better visible-thermal fusion method in the context of on-orbit inspection.